OpenStudy (anonymous):

solve the system of equations algebraically 1/3x+1/3y=1 and 2x+2y=9

OpenStudy (phi):

is the 1st equation \[ \frac{1}{3}x + \frac{1}{3}y = 1 \text{ ?} \]

OpenStudy (anonymous):

yes

OpenStudy (phi):

if so, multiply each term by 3 (multiply both sides of the equation by 3) to get x +y =3 now you have x+y=3 2x+2y=9 or x+y=3 x+y = 9/2 (divided both sides by 2) What 2 numbers, when added together equal 3 and at the same time equal 9/2 ? None. There is no solution for this pair of equations